OpenStudy (loser66):

a^2+b^2+c^2+ab+bc+ac>0 for all a, b, c? please, help

OpenStudy (loser66):

I need prove or disprove the expression

OpenStudy (aravindg):

You can try relating it to (a+b+c)^2 by adding and subtracting terms.

OpenStudy (loser66):

I did, but go nowhere (a+b+c)^2 -ab-bc-ac and then?

OpenStudy (aravindg):

Try this then (a-b)^2 + (b-c)^2 + (a-c)^2 >= 0

OpenStudy (aravindg):

We know that this is always true. Expand and reach your inequality.

ganeshie8 (ganeshie8):

\(\large a^2 + b^2 + c^2 +ab+bc+ca = \frac{1}{2}[(a+b)^2 +(b+c)^2 + (c+a)^2 ]\)

ganeshie8 (ganeshie8):

which is > 0, if \(a\ne -b, b\ne -c, c\ne -a\)

ganeshie8 (ganeshie8):

u can compress it into a single constraint

OpenStudy (loser66):

Thank you. I know what the result is now. So, it is not >0 for all. counter example is a = 1, b= -1 c =-1?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

try a = 1, b = -1, c = 1

ganeshie8 (ganeshie8):

ahh doesnt work either

OpenStudy (loser66):

if it >0 , need prove, if it is not >0, just counter example. That's all @AravindG I expand your expression, it goes to something different from original one :)

ganeshie8 (ganeshie8):

your counter example doesnt seem to work... check once

OpenStudy (loser66):

yes, it doesn't hehehe

ganeshie8 (ganeshie8):

counter example : a = 0, b = 0, c = 0

ganeshie8 (ganeshie8):

not an exciting one.. .

OpenStudy (aravindg):

I misread the question and tried to prove a^2 + b^2 + c^2 ≥ ab + ac + bc. Sorry.

OpenStudy (loser66):

but if a = b=0, c =-1 works

ganeshie8 (ganeshie8):

doesnt work

OpenStudy (loser66):

again, you are right, me wrong, hehehehe

OpenStudy (kinggeorge):

If we just use the expansion kindly provided by @ganeshie8, and the fact that it's equal to 0 if and only if \(a=-b, b= -c, c=-a\), then we can solve the expression in terms of \(a\). We get\[b=-a\]\[c=-b=a\]but also\[c=-a\]So \(c=a=-a\). This is only true if \(a=0\), and consequently \(a=b=c=0\). So other than the trivial solution where they're all 0, your expression must always be greater than 0.

ganeshie8 (ganeshie8):

wow ! that looks really cool+simple proof xD

OpenStudy (loser66):

Thank you very much, I got it.